OpenStudy (anonymous):

Really need these answers! Evaluate each of the following without using a calculator or computer. Show your work. a. sin(arcsin(1/2) + arcos(1/2)) b. sec(arccos(-√3/2)) c. arctan(tan(3π/4)) d. tan(arctan(3π/4))

OpenStudy (gorv):

sin(arcsin(1/2) + arcos(1/2)) arcos(1/2)=pi/3 arcsin(1/2)=pi/6 pi/3+pi/6=pi/2 sin pi/2=1

OpenStudy (gorv):

arccos(-√3/2)= pi-arccos(√3/2)=pi-pi/6=2*pi/3 sec(2*pi/3)=\[-\sqrt{3}/2\]

OpenStudy (gorv):

c tan inverse cancels tan ans=3*pi/4

OpenStudy (gorv):

similarly for d